OpenStudy (dallaslifebaby01):

y = -2.8x + 150.4 y = 14.4x + 102.8 can someone solve and show steps? i already did i just want to check my work

OpenStudy (dallaslifebaby01):

my solution -2.8x+150.4 = 14.4x +102.8 150.4 = 14.4x +2.8 + 102.8 150.4 – 102.8 = 17.2x 47.6 = 17.2x x=47.6/17.2

OpenStudy (amorfide):

-2.8x+150.4 = 14.4x +102.8 correct 150.4 = 14.4x +2.8 + 102.8 you forgot to put x after the 2.8, so this line should be 150.4 = 14.4x +2.8x + 102.8 150.4 – 102.8 = 17.2x correct 47.6 = 17.2x correct x=47.6/17.2 correct

OpenStudy (dallaslifebaby01):

is that all i would need to solve? I'm not missing anything? and its completely solved?

OpenStudy (amorfide):

you cuold probably use a calculator to get a decimal answer

OpenStudy (amorfide):

you would then use that value of x, substitute it into one of the two given equations and work out a y value

OpenStudy (dallaslifebaby01):

okay thanks

OpenStudy (amorfide):

y = 14.4x + 102.8 y=14.4(47.6/17.2)+102.8

OpenStudy (amorfide):

then you have both values then it is solved

OpenStudy (dallaslifebaby01):

okay thanks